Use substitution y = 2^X to solve for X the equation 2^(2X +1)– 2^(X +1) + 1 = 2X


1
Expert's answer
2021-09-14T11:34:37-0400

Let "y=2^x, y>0"

Substitute


"2y^2-2y+1=y"

"2y^2-3y+1=0"

"y(2y-1)-(2y-1)=0"

"(2y-1)(y-1)=0"

"y_1=\dfrac{1}{2}, y_2=1"

"2^x=\dfrac{1}{2}"

"x=-1"

"2^x=1"

"x=0"

"x_1=-1, x_2=0"


"\{-1, 0\}"